OpenStudy (anonymous):

The equation of the line that goes through the point ( 9 ,8 ) and is perpendicular to the line 4 x + 2 y = 2 can be written in the form y = mx+b where m is: _____ and where b is: ____ Please, help!!!

OpenStudy (anonymous):

m = 1/2 *0.5* b = 7/2 *3.5*

OpenStudy (anonymous):

is it right???

OpenStudy (anonymous):

Yes, thank you!!

OpenStudy (anonymous):

do u know how i got it thou???